# Benchmark brown dwarfs – I. A blue M2 + T5 wide binary and a probable young M4 + [T7 + T8] hierarchical triple

### v1

@@ 5 Summary and conclusions
We have conducted a search for late T dwarf wide binaries using CatWISE and Gaia DR3, resulting in the discovery of two new M + T dwarf wide binary systems. Both systems are confirmed to be gravitationally bound and stable based on their tidal radii. Optical and NIR spectroscopic follow-up observations were performed for all companions, enabling spectral classification and characterization.
The first system, L 122-88 AB, exhibits spectral features indicative of mild metal-poor composition, including a weaker 710 nm TiO absorption band in L 122-88 A and suppressed K-band flux in L 122-88 B. The best-fitting BT-Settl model to the 0.3-2.5 m spectrum of L 122-88 A yields = 3500 K, [Fe/H] = , and log = 5.0. We classify L 122-88 AB as a mildly metal-poor M2 + T5 wide binary.
The second system, UPM J10403551 AB, is likely a hierarchical triple system consisting of an unresolved T7 + T8 binary and an M4 dwarf wide companion at a distance of 25 pc from the Sun. While UPM J10403551 Ba/Bb are unresolved in survey images, their NIR spectrum is best fitted by a combined T7 + T8 template, explaining the observed excess luminosity. This system might be the third known triple system similar to Indi ABab or Gliese 229 ABab. Based on the H emission of the M4 primary, we estimate the system age to be 0.3-2 Gyr. At this relatively young age, the T8 component UPM J10403551 Bb has a mass between 9 and 28 Jupiter masses and potentially is a planetary mass object.
@@ Acknowledgements
Based on observations obtained at the Southern Astrophysical Research (SOAR) telescope, which is a joint project of the Ministério da Ciência, Tecnologia e Inovações (MCTI/LNA) do Brasil, the US National Science Foundation’s NOIRLab, the University of North Carolina at Chapel Hill (UNC), and Michigan State University (MSU). Based on observations collected at the European Southern Observatory under ESO programme 113.26V7.001.
@@
This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This publication makes use of data products from the Wide field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Lab- oratory/California Institute of Technology, funded by the National Aeronautics and Space Administration This publication makes use of data products from the Wide-field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. SPLAT is an collaborative project of research students in the UCSD Cool Star Lab, aimed at developing research through the building of spectral analysis tools. Contributors to SPLAT have included Christian Aganze, Jessica Birky, Daniella Bardalez Gagliuffi, Adam Burgasser (PI), Caleb Choban, Andrew Davis, Ivanna Escala, Joshua Hazlett, Carolina Herrara Hernandez, Elizabeth Moreno Hilario, Aishwarya Iyer, Yuhui Jin, Mike Lopez, Dorsa Majidi, Diego Octavio Talavera Maya, Alex Mendez, Gretel Mercado, Niana Mohammed, Johnny Parra, Maitrayee Sahi, Adrian Suarez, Melisa Tallis, Tomoki Tamiya, Chris Theissen, and Russell van Linge.
@@
Z.H.Z acknowledges the supports from the Jiangsu Province fundamental research programme (BK20211143), the Program for Innovative Talents and Entrepreneur in Jiangsu (JSSCTD202139), and the science research grants from the China Manned Space Project with NO. CMS-CSST-2021-A08. The work of F.N. is supported by NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. M.C.G.O. acknowledges financial support from the Agencia Estatal de Investigación (AEI/10.13039/501100011033) of the Ministerio de Ciencia e Innovación and the ERDF “A way of making Europe” through project PID2022-137241NB-C42. N.L. acknowledges support from the Agencia Estatal de Investigación del Ministerio de Ciencia e Innovación (AEI-MCINN) under grant PID2022-137241NB-C41. B.G. acknowledges support from the Polish National Science Center (NCN) under SONATA grant No. 2021/43/D/ST9/0194. P.C. acknowledges financial support from the Spanish Virtual Observatory project funded by the Spanish Ministry of Science and Innovation/State Agency of Research MCIN/AEI/10.13039/501100011033 through grant PID2020-112949GB-I00. The authors would like to thank the anonymous reviewer for his/her valuable comments.

### v3

@@ 5 Summary and conclusions
We have conducted a search for late T dwarf wide binaries using CatWISE and Gaia DR3, resulting in the discovery of a wide binary and a probable hierarchical quadruple system. Both systems are confirmed to be gravitationally bound and stable based on their tidal radii. Optical and NIR spectroscopic follow-up observations were performed for all companions, enabling spectral classification and characterization.
The first system, L 122-88 AB, exhibits spectral features indicative of mildly metal-poor composition, including a weaker 710 nm TiO absorption band in L 122-88 A and suppressed -band flux in L 122-88 B. The best-fitting BT-Settl model to the 0.3-2.5 m spectrum of L 122-88 A yields = 3500 K, [Fe/H] = , and log = 5.0. We classify L 122-88 AB as a mildly metal-poor M2 + T5 wide binary.
The second system, UPM J10403551 AB, is likely a hierarchical quadruple system consisting of a T7 + T8 spectral binary and an M4 + M4 astrometric binary at a distance of 25 pc from the Sun. Both the RUWE (1.47) of UPM J10403551 A and its excess luminosity (0.7 mag brighter than model predications of its ) suggest it is an unresolved M4 + M4 binary. While UPM J10403551 Ba/Bb are unresolved in survey images, their NIR spectrum is best fitted by a combined T7 + T8 template, explaining the observed excess luminosity. UPM J10403551 AabBab is the first known quadruple system with a close binary of T dwarfs. Based on the H emission of the M4 primary, we estimate the system age to be 0.3-2 Gyr. At this relatively young age, the T8 component UPM J10403551 Bb has a mass between 9 and 28 Jupiter masses and potentially is a planetary mass object.
@@ Acknowledgements
This study is based on observations obtained at the SOAR telescope, which is a joint project of the Ministério da Ciência, Tecnologia e Inovações (MCTI/LNA) do Brasil, the US National Science Foundation’s NOIRLab, the University of North Carolina at Chapel Hill (UNC), and Michigan State University (MSU). This study is based on observations collected at the European Southern Observatory under ESO programme 113.26V7.001.
@@
This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This publication makes use of data products from the WISE, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. SPLAT is a collaborative project of research students in the UCSD Cool Star Lab, aimed at developing research through the building of spectral analysis tools. Contributors to SPLAT have included Christian Aganze, Jessica Birky, Daniella Bardalez Gagliuffi, Adam Burgasser (PI), Caleb Choban, Andrew Davis, Ivanna Escala, Joshua Hazlett, Carolina Herrara Hernandez, Elizabeth Moreno Hilario, Aishwarya Iyer, Yuhui Jin, Mike Lopez, Dorsa Majidi, Diego Octavio Talavera Maya, Alex Mendez, Gretel Mercado, Niana Mohammed, Johnny Parra, Maitrayee Sahi, Adrian Suarez, Melisa Tallis, Tomoki Tamiya, Chris Theissen, and Russell van Linge.
@@
ZHZ acknowledges the supports from the Jiangsu Province fundamental research programme (BK20211143), the Program for Innovative Talents and Entrepreneur in Jiangsu (JSSCTD202139), and the science research grants from the China Manned Space Project with no. CMS-CSST-2021-A08. The work of FN is supported by NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. MCGO acknowledges financial support from the Agencia Estatal de Investigación (AEI/10.13039/501100011033) of the Ministerio de Ciencia e Innovación and the ERDF ’A way of making Europe’ through project PID2022-137241NB-C42. NL acknowledges support from the Agencia Estatal de Investigación del Ministerio de Ciencia e Innovación (AEI-MCINN) under grant PID2022-137241NB-C41. BG acknowledges support from the Polish National Science Center (NCN) under SONATA grant no. 2021/43/D/ST9/0194. PC acknowledges financial support from the Spanish Virtual Observatory project funded by the Spanish Ministry of Science and Innovation/State Agency of Research MCIN/AEI/10.13039/501100011033 through grant PID2020-112949GB-I00. The authors would like to thank the anonymous reviewer for his/her valuable comments.